OpenStudy (decarr432):

if f(x)=3x+3, then f^-1(x)=?

OpenStudy (astrophysics):

To find the inverse: Replace f(x) with y Switch x's and y's, so put x where y is and y where x is. Solve for y Replace y with f^-1(x)

OpenStudy (decarr432):

Huh?

OpenStudy (astrophysics):

Those are the steps for solving the inverse

OpenStudy (decarr432):

Okay

OpenStudy (astrophysics):

So lets go over this step by step, what does it say we should do first

OpenStudy (decarr432):

Replace f(x) with the y

OpenStudy (astrophysics):

Yes, so \[y = 3x+3\] next?

OpenStudy (decarr432):

Switch the x's and y's

OpenStudy (astrophysics):

Right interchange them

OpenStudy (decarr432):

wouldn't it be 3x+3 still?

OpenStudy (astrophysics):

\[x=3y+3\]

OpenStudy (astrophysics):

Now simply solve for y

OpenStudy (decarr432):

Okay how so? I don't understand how

OpenStudy (astrophysics):

We need to use algebra, remember how I mentioned how important it is :P Ok so since we are solving for y, lets first get rid of the +3 on the right side, what would we do?

OpenStudy (decarr432):

subtract 3 from both sides

OpenStudy (astrophysics):

Exactly

OpenStudy (astrophysics):

\[x-3=3y\] now we need to get rid of that 3 being multiplied to the y, how can we move it?

OpenStudy (decarr432):

divide by 3?

OpenStudy (astrophysics):

Right!

OpenStudy (astrophysics):

\[y = \frac{ x-3 }{ 3 } = \frac{ x }{ 3 }-1\]

OpenStudy (astrophysics):

Good?

OpenStudy (decarr432):

yeah

OpenStudy (astrophysics):

Ok now last step is to replace y with f^(-1)

OpenStudy (astrophysics):

And behold we have our inverse!

OpenStudy (decarr432):

Okay so can you help me set it up?

OpenStudy (decarr432):

and sorry if i'm keeping you up its late

OpenStudy (astrophysics):

It's done, that's our answer \[y= \frac{ x }{ 3 }-1 \implies f^{-1}(x) = \frac{ x }{ 3 }-1\]

OpenStudy (decarr432):

okay

OpenStudy (decarr432):

thanks bud

OpenStudy (decarr432):

I'm done for the night have a good one

OpenStudy (astrophysics):

Yw :), take care

OpenStudy (astrophysics):

Here is a video on explaining what exactly this means, I recommend you watch it: https://www.khanacademy.org/math/algebra2/manipulating-functions/introduction-to-inverses-of-functions/v/introduction-to-function-inverses